I'm trying to figure it out but I'm not coming up with the right answer. According to my book, the answer should be 11, 13, 15...HELP Please! Mike
2X+4+7 = X+2 X+4

Answer by kunal_122(19)

the answer in your textbook is right.
i think thats the equation
2x-7=x+4
2x-x=7+4
x=11
x+2=13 and x+4=15 and u can check the answer smallest interger is x and its 11 . twic the 11 is 22 which is 7 more larger than 15 the largest
2